traveling 50 mph how long would it take to travel 550 miles ?

traveling 60 mph how long would it take to travel 660 miles?
traveling 80 mph how long would it take to travel 320 miles?

Answer by jessica43(140)

You need to rewrite these as formulas and then you can solve them:
Traveling 50 mph how long would it take to travel 550 miles
50*H = 550 (where H = the number of hours)
H = 11 hours
traveling 60 mph how long would it take to travel 660 miles
60*H = 660
H = 11 hours
traveling 80 mph how long would it take to travel 320 miles
80*H = 320
H = 4 hours